OpenStudy (anonymous):

You are installing a fence at the City Park. The square field must be 484 square feet. If the fence cost $1.75 per foot, how much will the fence cost?

OpenStudy (anonymous):

Since we know that a square has equal sides we can use LxW=A Since L and W would be the same (square has equal sides) we can set it as L^2. So the equation we have is 484 = L^2. So if we take the root of 484 we can solve for L. Therefore L = 22 Since the fence costs $1.75 we multiply 1.75 * 22 giving us 38.5. So the fencing would cost $38.5

OpenStudy (anonymous):

Because it is a square field, to find the side , you need to square root the area, \[\sqrt{484}=22\] Then, find the perimeter of the field \[22*4=88\] Finally, Cost = 88X1.75=154

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

yall giving two different answers

OpenStudy (anonymous):

Im pretty sure mines right. Im in grade 12 calculus

OpenStudy (anonymous):

Im pretty sure mines right. I am in grade 13 calculus